OpenStudy (anonymous):

When solving the system of equations by graphing, what is the solution of 3x + 2y = 2 and 2x – y = 6? A. (–2, 2) B. (2, –2) C. (–2, –2) D. (2, 2)

OpenStudy (anonymous):

3x+2y=2 in slope intercept form is y=-3/2x+1 2x-y=6 in slope intercept form is y= 2x-6 now if u graph those u will get the solution